You pour 210 g hot coffee at 78.7 degree C and som
Answer #1

The first problem have lost information.

22.- the energy added or removed from a material with a temperature change is calculated as

\small Q = mc\Delta T

where \small m is the mass of material, \small c the specific heat, and \small \Delta T the temperature change during the energy change

the zero law of thermodynamics stablidhes that the energy flow is from hot objects to cold objects, and the conservation of energy say us,

\small Q_{cold} = -Q_{hot}

the energy lost by the hot object is gained by the cold object.

In this case the coffe lost energy and the creap and cup gain it. that is

\small -Q_{coffe} = Q_{cream} + Q_{cup}

\small -m_{coffe}c_{coffe}\Delta T_{coffe} = m_{cream}c_{cream}\Delta T_{cream} + m_{cup}c_{cup}\Delta T_{cup}

We know all that variables except the mass of the cream, then solving for that we have

\small m_{cream} = \frac{-m_{coffe}c_{coffe}\Delta T_{coffe} - m_{cup}c_{cup}\Delta T_{cup}}{c_{cream}\Delta T_{cream}}

\small m_{cream} = \frac{-\left ( 0.210\textup{ kg} \right )\left ( 1.00\textup{ kcal/kg}^\circ \textup{C} \right )\left ( 63^\circ \textup{C} - 78.7^\circ \textup{C} \right ) - \left ( 0.115\textup{ kg} \right )\left ( 0.260\textup{ kcal/kg}^\circ \textup{C} \right )\left ( 63^\circ \textup{C} - 22.0 ^\circ \textup{C} \right )}{1.00 \textup{kcal/kg} ^\circ \textup{C}\left (63^\circ \textup{C} - 7.50^\circ \textup{C} \right )}

\small m_{cream} = \frac{3.297\textup{ kcal} - 1.2259\textup{ kcal}}{55.5 \textup{kcal/kg}}

\small m_{cream} = 0.0373\textup{ kg}

m_{cream} = 37.3\textup{ g}

23.- The heat required to raise the temperature of water from 0°C to 25°C is

Q = mc\Delta T

Q = \left (0.740\textup{ kg} \right )\left ( 1.00\textup{ kcal/kg} ^\circ\textup{C} \right )\left ( 25^\circ \textup{C} - 0^\circ \textup{C}\right )

Q = 18.5\textup{ kcal}